I need Emergency help solving this problem!! Can someone please respond!!
The Everton College store paid $1848 for an order of $51 calculations. The store ore paid $8 for each scientific calculator. The others, all graphing calculators cost the store $53. How many of each type of calculator was ordered? The store order____ scientific calculator and____ graphing calculator___

Answer by stanbon(75887)

The Everton College store paid $1848 for an order of 51 calculations. The store ore paid $8 for each scientific calculator. The others, all graphing calculators cost the store $53. How many of each type of calculator was ordered?
-----------
Equations:
Quantity: s + g = 51
Value::: 8s +53g = 1848
--------------------
Multiply thru Quantity by 8.
-----
8s + 8g = 8*51
8s +53g = 1848
-----
Subtract and solve for "g":
45g = 1440
g = 32 (# of graphing calculators bought)
-------------
solve for "s":
s + g = 51
s = 51-32 = 19 (# of scientific calculators bought)
=============
Cheers,
Stan H.
=============





The store order____ scientific calculator and____ graphing calculator___